# Transfer Matrix Approach for Topological Edge States

### v2

@@ IV.1 Tetramer SSH4 Model
Recently there has been considerable interest, in both theory and experiment, in increasing the number of sites per unit cell [27, 28, 13, 29, 30]. In the trimer case, the results are not too interesting. Numerical calculations show that the number of edge states simply correspond to the number of sites to the left of the weakest coupling. The three possible cuts are shown in Fig. 3, and we see that the number of edge modes agrees with our intuition.
@@
Figure 3: Summary of the numerical results for the three different cuts in the SSH3 model. The number of left edge modes are given on the left, and they correspond to the number of sites to the left of the weakest coupling (modulo 3). This agrees with the grouping argument frequently used for one-dimensional topological systems [31].
Adding another site per unit cell, numerical calculations for the SSH4 model show that this intuitive picture does not always hold. As such, it may prove useful to investigate analytically what happens when we have $4$ sites per unit cell.
@@
$T=\left(\begin{array}[]{cc}-\frac{\gamma_{2}}{\gamma_{3}}&-\frac{\mu_{3}}{\gamma_{3}}\\ \frac{\gamma_{2}\mu_{4}}{\gamma_{3}\gamma_{4}}&\frac{\mu_{3}\mu_{4}}{\gamma_{3}\gamma_{4}}-\frac{\gamma_{3}}{\gamma_{4}}\\ \end{array}\right)\left(\begin{array}[]{cc}-\frac{\gamma_{4}}{\gamma_{1}}&-\frac{\mu_{1}}{\gamma_{1}}\\ \frac{\gamma_{4}\mu_{2}}{\gamma_{1}\gamma_{2}}&\frac{\mu_{1}\mu_{2}}{\gamma_{1}\gamma_{2}}-\frac{\gamma_{1}}{\gamma_{2}}\\ \end{array}\right)$ (16)
@@
We see that when $\bar{\mu}=0$, the $T$ matrix may support a topological edge state with $T$-eigenvalue $\gamma_{3}\gamma_{1}/(\gamma_{4}\gamma_{2})$. That is, the SSH4 model seems to act just like 2 jagged copies of the dimer SSH model, where the $T$-eigenvalue for the edge state is just the product of the eigenvalues for each dimer. However, we note that, even if $\gamma_{1}>\gamma_{2}$, the SSH4 chain may support a topological edge state as long as $\gamma^{2}_{13}/\gamma_{2}<\gamma_{4}$. Furthermore, even if $\gamma_{4}$ is the weakest coupling, we can always choose $\gamma_{13},\gamma_{2}$ so that the eigenvalue remains less than 1, and an edge state is supported. This shows that the ‘limiting to different groupings’ argument frequently used in the discussion of the SSH model and the Kitaev chain (see Refs. [31, 32] and Fig. 3) is not so easily generalisable to more complicated systems. This is not the only topological edge state of the SSH4 model. The $T$ matrix also supports edge states when $\bar{\mu}=\pm\sqrt{\gamma_{13}^{2}+\gamma_{2}^{2}}$. In this case the eigenvalue is $\gamma_{2}/\gamma_{4}$, and it looks like the edge state of an SSH chain with the $\gamma_{1}=\gamma_{3}$ couplings ignored.
@@
Computing the Zak phase of each band as in Fig. 4 shows that it is quantised. This verifies that our restricted SSH4 model is indeed topologically nontrivial. Our analysis also shows that the group of topological invariants for the SSH4 model is $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
So far we have only analysed the SSH4 model protected by parity symmetry. In the dimer case, both parity symmetry and chiral symmetry impose the same requirement $v_{1}=v_{2}$, and both are individually responsible for the topological protection of the edge mode [25]. However, in the tetramer SSH4 case chiral symmetry imposes a different condition $v_{i}=0$ than parity symmetry. For a brief analysis, on the transfer matrix level, chiral symmetry is satisfied as long as
$\sigma_{z}T(E)\sigma_{z}=T(-E).$
@@
Finally, our analysis of the SSH4 model shows that unlike the Bloch Hamiltonian, the transfer matrix lends itself easily to expanding the size of the unit cell. This comes at the expense of making long-range couplings more difficult to introduce. However, this also means that the Hamiltonians of continuous systems can still be described by $2\times 2$ transfer matrices, and as such is susceptible to our analysis thus far. Our method therefore generalises to any real one-dimensional quantum mechanical or photonic systems, without discretisation.
@@ IV.2 Dimerised Photonic Crystal
The transfer matrix is a well-known tool in analysing photonic crystals, and have previously been used to prove a bulk-edge correspondence in the interface of two inversion symmetric one-dimensional photonic crystals [26].
@@
$\hat{T}_{\Lambda}=\prod_{i=N}^{1}\begin{pmatrix}\cos(\phi_{i})&\mathrm{i}\mkern 1.0muZ_{i}^{-1}\sin(\phi_{i})\\ \mathrm{i}\mkern 1.0muZ_{i}\sin(\phi_{i})&\cos(\phi_{i})\end{pmatrix}$ (20)
where $\phi_{i}=k_{x,i}h_{i}$, $Z_{i}^{TE}=Z_{0}\mu_{i}k_{0}/k_{x,i}$ in the TE case, and $Z_{i}^{TM}=Z_{0}k_{x,i}/(\varepsilon_{i}k_{0})$ in the TM case. Here, $Z_{0}$ is the impedance of free space, $k_{0}=\sqrt{\varepsilon_{i}\mu_{i}}\omega/c$ is the total wavenumber, so that $k_{x,i}=\sqrt{\varepsilon_{i}\mu_{i}k_{0}^{2}-\beta^{2}}$ where $\beta$ is the wavenumber in the continuous-translation-invariant direction ($z$ direction in Fig. 5 (a)).
@@
Then, by changing basis to
@@
the impedance matching condition on the left and right edges become $F_{-}=0$ and $F_{+}=0$ respectively. As such, the condition for an edge mode becomes identical to the one we obtained for the SSH models:
$\bar{T}_{\Lambda}(\beta)\begin{pmatrix}0\\ 1\end{pmatrix}=a\begin{pmatrix}0\\ 1\end{pmatrix}{,\ \ \ }a\in(-1,1).$ (23)
@@
We may compute the band structure of TM waves in the DPC with $q=1$ using the $T$-matrix numerically, with the bulk bands computed using Bloch’s theorem, and the edge states by numerically finding solutions to (23) in the following manner: The parameter space spanned by $\beta$ and $\delta h$ is discretised to a grid of size $500\times 50$ respectively. A cell is then marked as an edge state if (23) is satisfied within an adjustable tolerance. This results Fig. 5 (b) and we see that the results are unlike the SSH model. The edge state appears on the wrong side $\delta h<0$, which in the coupled mode approximation (Appendix B) means $\gamma_{1}>\gamma_{2}$. In addition, we see an extended edge state appearing below the two bands, which spans through both regions $\delta h<0$ and $\delta h>0$.
The condition that $T$ is parity invariant, Eq. (7), is exactly the condition that the unit cell is symmetric $q=1/2$. In this case, all edge states are topological and the $\mathrm{SL}(2,\mathbb{R})$ trajectories for the TM case are illustrated in Fig. 6. The TE case gives similar results, so we will omit it for brevity. From this figure, we see that edge states are supported if and only if $\delta h>0$, in agreement with the coupled mode approximation in Appendix B. There seems to be no extended edge states in this case. As a result, we see that the extended edge mode in Fig. 5 (b) is a result of the defect introduced by the asymmetric prescription of the unit cell, and is not a property of the bulk.
@@
We have characterised the edge states of the DPC for the allowed values of $\beta$ such that the field decays away from the periodic structure. The extended edge state is understood to arise from the defect introduced by the asymmetric perturbation $q\neq 0$, and its existence seems to be weakly tied to the topological edge state between the two SSH-like bands predicted by the coupled mode approximation (Appendix B).
@@ Appendix A Proof of Zak Phase-Eigenvector Phase relation
Consider a system with $N$ sites per unit cell. Without loss of generality, we may assume that $N$ is even. We label the amplitude in the $n^{\text{th}}$ site as $a_{n}$. In general, due to $PT$ symmetry, the Bloch eigenvector will satisfy
@@
$\left|K\right\rangle=\frac{\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}}{\sqrt{1+r_{2}^{2}(k)+\cdots}}\begin{pmatrix}\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu\phi(K)/2}\\ r_{2}(K)\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu\phi_{2}(K)}\\ \vdots\\ r_{2}(K)\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi_{2}(K)}\\ \mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}\end{pmatrix}.$ (25)
Note that the choice of reciprocal gauge $\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}$ is physical: It is required for the Zak phase to have an unambiguous interpretation of mean displacement. To prove this, we first fix the first site to have position $x_{1}=0$, so that the lattice position of the $n^{\text{th}}$ site is given by $x_{n}=n-1$. As a result, the Bloch eigenvector is related to the discrete fourier transform (DFT) by
@@
$\langle n=1|K,\text{Bloch}\rangle=\langle n=1|K,\text{DFT}\rangle=a_{1}=\text{const.}$. (28)
We may fix the global phase to 00, so that the Bloch eigenvector is given by (25). We thus see that allowing the phase of the first entry in $\left|K\right\rangle$ to vary is equivalent to choosing a different, translated coordinate system for each value of the crystal momentum $K$.
@@
$=\left\langle K,\text{Sym}\right|\left[-\mathrm{i}\mkern 1.0mu\partial_{K}+\frac{\phi^{\prime}(K)}{2}\right]\left|K,\text{Sym}\right\rangle$ (29)
While computing the inner product $\left\langle K,\text{Sym}\right|\partial_{K}\left|K,\text{Sym}\right\rangle$, the $k^{\text{th}}$ and $(N-k)^{\text{th}}$ terms always cancel, so we are left with
@@
The generalisation to continuous systems is straightforward by taking $N\to\infty$.
Appendix B Coupled Mode Approximation to the Dimerised Photonic Crystal

### v3

@@ IV.1 Tetramer SSH4 Model
Recently there has been considerable interest, in both theory and experiment, in increasing the number of sites per unit cell [31, 32, 14, 33, 34, 35, 36]. In the trimer case, the results are not too interesting. Numerical calculations show that the number of edge states simply correspond to the number of sites to the left of the weakest coupling. The three possible cuts are shown in Fig. 4, and we see that the number of edge modes agrees with our intuition.
@@
Figure 4: Summary of the numerical results for the three different cuts in the SSH3 model. We may interpret the bulk modes as modes that arise due to the collective excitation of the red-boxed sites to the right of the weakest coupling (single-lines). The edge modes are then interpreted as the excitation of blue-boxed leftover sites. The number of left edge modes are given on the left, and they correspond to the number of sites to the left of the weakest coupling (modulo 3). This agrees with the grouping argument frequently used for one-dimensional topological systems [37].
Adding another site per unit cell, numerical calculations for the SSH4 model show that this intuitive picture does not always hold. As such, it may prove useful to investigate analytically what happens when we have $4$ sites per unit cell. Explicitly, this SSH4 model is described by the following Hamiltonian, analogous to equation (II):
$H=$ $\sum\limits_{m}\gamma_{1}\left|m,2\right\rangle\left\langle m,1\right|+\gamma_{2}\left|m,3\right\rangle\left\langle m,2\right|$
$+\gamma_{3}\left|m,4\right\rangle\left\langle m,3\right|+\gamma_{4}\left|m+1,1\right\rangle\left\langle m,4\right|+\mathrm{h.c.}$
$+v_{1}\left|m,1\right\rangle\left\langle m,1\right|+v_{2}\left|m,2\right\rangle\left\langle m,2\right|$
$+v_{3}\left|m,3\right\rangle\left\langle m,4\right|+v_{4}\left|m,4\right\rangle\left\langle m,4\right|.$ (16)
@@
$T=\left(\begin{array}[]{cc}-\frac{\gamma_{2}}{\gamma_{3}}&-\frac{\mu_{3}}{\gamma_{3}}\\ \frac{\gamma_{2}\mu_{4}}{\gamma_{3}\gamma_{4}}&\frac{\mu_{3}\mu_{4}}{\gamma_{3}\gamma_{4}}-\frac{\gamma_{3}}{\gamma_{4}}\\ \end{array}\right)\left(\begin{array}[]{cc}-\frac{\gamma_{4}}{\gamma_{1}}&-\frac{\mu_{1}}{\gamma_{1}}\\ \frac{\gamma_{4}\mu_{2}}{\gamma_{1}\gamma_{2}}&\frac{\mu_{1}\mu_{2}}{\gamma_{1}\gamma_{2}}-\frac{\gamma_{1}}{\gamma_{2}}\\ \end{array}\right).$ (21)
@@
We see that when $\bar{\mu}=0$, the $T$ matrix may support a topological edge state with $T$-eigenvalue $\gamma_{3}\gamma_{1}/(\gamma_{4}\gamma_{2})$. That is, the SSH4 model seems to act just like 2 jagged copies of the dimer SSH model, where the $T$-eigenvalue for the edge state is just the product of the eigenvalues for each dimer. However, we note that, even if $\gamma_{1}>\gamma_{2}$, the SSH4 chain may support a topological edge state as long as $\gamma^{2}_{13}/\gamma_{2}<\gamma_{4}$. Furthermore, even if $\gamma_{4}$ is the weakest coupling, we can always choose $\gamma_{13},\gamma_{2}$ so that the eigenvalue remains less than 1, and an edge state is supported. This shows that the ‘limiting to different groupings’ argument frequently used in the discussion of the SSH model and the Kitaev chain (see Refs. [37, 38] and Fig. 4) is not so easily generalisable to more complicated systems.
This is not the only topological edge state of the SSH4 model. The $T$ matrix also supports edge states when $\bar{\mu}=\pm\sqrt{\gamma_{13}^{2}+\gamma_{2}^{2}}$. In this case the eigenvalue is $\gamma_{2}/\gamma_{4}$, and it looks like the edge state of an SSH chain with the $\gamma_{1}=\gamma_{3}$ couplings ignored. Our analysis using the $T$-matrix thus shows that there are two quantities which are independently responsible for the creation and destruction of topological edge states. Namely, $\gamma_{13}^{2}/(\gamma_{2}\gamma_{4})$ and $\gamma_{2}/\gamma_{4}$.
@@
We will now verify that our $T$-matrix results are in agreement with the conventional method of computing Bloch-band topological invariants. We compute the Zak phase of each band using (12). The results, summarised in Fig. 5, show that the Zak phase is quantised to multiples of $\pi$. This quantisation of the Zak phase is evidence that the restricted SSH4 model is indeed topologically nontrivial. We note that there are two families of edge states which can be created or destroyed independently of the other. The first family consists of the edge state between [^4] As in figure 5 we number the bands from lowest to highest in energy. bands 2 and 3, and the second family consists of the edge states between bands 1 and 2, and between bands 3 and 4. Each family can only support 1 or 0 edge states. As such, the topological phases of the SSH4 model can be classified by the group is $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
So far we have only analysed the SSH4 model protected by parity symmetry. In the dimer case, both parity symmetry and chiral symmetry [21, Sec. II.C] impose the same requirement $v_{1}=v_{2}$ in Eq. (II), and both are individually responsible for the topological protection of the edge mode [20]. However, in the tetramer SSH4 case chiral symmetry imposes a different condition $v_{i}=0$ than parity symmetry, see Eq. (IV.1). For a brief analysis, on the transfer matrix level, chiral symmetry is satisfied as long as
$\sigma_{z}T(E)\sigma_{z}=T(-E).$ (25)
@@
[…]
@@ IV.2 Dimerised Photonic Crystal
The transfer matrix is a well-known tool in analysing photonic crystals, and have previously been used to prove a bulk-edge correspondence in the interface of two inversion symmetric one-dimensional photonic crystals [30]. We note that our transfer matrix approach is especially useful for photonic systems since the analogous Hamiltonian for photonic systems is more difficult to work with than the transfer matrix [20].
@@
$\hat{T}_{\Lambda}=\prod_{i=N}^{1}\begin{pmatrix}\cos(\phi_{i})&\mathrm{i}\mkern 1.0muZ_{i}^{-1}\sin(\phi_{i})\\ \mathrm{i}\mkern 1.0muZ_{i}\sin(\phi_{i})&\cos(\phi_{i})\end{pmatrix},$ (26)
where $\phi_{i}=k_{x,i}h_{i}$, $Z_{i}^{TE}=Z_{0}\mu_{i}k_{0}/k_{x,i}$ in the TE case, and $Z_{i}^{TM}=Z_{0}k_{x,i}/(\varepsilon_{i}k_{0})$ in the TM case. Here, $Z_{0}$ is the impedance of free space, $k_{0}=\sqrt{\varepsilon_{i}\mu_{i}}\omega/c$ is the total wavenumber, so that $k_{x,i}=\sqrt{\varepsilon_{i}\mu_{i}k_{0}^{2}-\beta^{2}}$ where $\beta$ is the wavenumber in the continuous-translation-invariant direction ($z$ direction in Fig. 6 (a)). The role of this transfer matrix is to take in the electroamgnetic waves at a certain position $x$ and calculate the electromagnetic waves one unit-cell away $x+\Lambda$. Namely, $\hat{T}_{\Lambda}:(E_{\perp}(x),H_{\perp}(x))\mapsto(E_{\perp}(x+\Lambda),H_{\perp}(x+\Lambda))$. Here, $\Lambda=\sum_{i}h_{i}$ is the unit cell length.
@@
Then, by changing basis from $(E_{\perp},H_{\perp})$ to $(F_{-},F_{+})$:
@@
the impedance matching condition (27) on the left and right edges become $F_{-}=0$ and $F_{+}=0$ respectively. As such, the condition for an edge mode becomes identical to the one we obtained for the SSH models:
$\bar{T}_{\Lambda}(\beta)\begin{pmatrix}0\\ 1\end{pmatrix}=\alpha\begin{pmatrix}0\\ 1\end{pmatrix}{,\ \ \ }\alpha\in(-1,1).$ (29)
@@
We may compute the band structure of TM waves in the DPC with $q=1$ using the $T$-matrix numerically, with the bulk bands computed using Bloch’s theorem, and the edge states by numerically finding solutions to (29) in the following manner: The parameter space spanned by $\beta$ and $\delta h$ is discretised to a grid of size $500\times 50$ respectively. A cell is then marked as an edge state if (29) is satisfied within an adjustable tolerance. This results Fig. 6 (b) and we see that the results are unlike the SSH model. The edge state appears on the wrong side $\delta h<0$, which in the coupled mode approximation (Appendix C) means $\gamma_{1}>\gamma_{2}$. In addition, we see an extended edge state appearing below the two bands, which spans through both regions $\delta h<0$ and $\delta h>0$.
The condition that $T$ is parity invariant, Eq. (10), is exactly the condition that the unit cell is symmetric $q=1/2$. In this case, all edge states are topological and the $\mathrm{SL}(2,\mathbb{R})$ trajectories for the TM case are illustrated in Fig. 7. The TE case gives similar results, so we will omit it for brevity. From this figure, we see that edge states are supported if and only if $\delta h>0$, in agreement with the coupled mode approximation in Appendix C. There seems to be no extended edge states in this case. As a result, we see that the extended edge mode in Fig. 6 (b) is a result of the defect introduced by the asymmetric prescription of the unit cell, and is not a property of the bulk.
@@
We have characterised the edge states of the DPC for the allowed values of $\beta$ such that the field decays away from the periodic structure. The extended edge state is understood to arise from the defect introduced by the asymmetric perturbation $q\neq 0$, and its existence seems to be weakly tied to the topological edge state between the two SSH-like bands predicted by the coupled mode approximation (Appendix C).
@@ Appendix A Proof of Zak Phase-Eigenvector Phase relation
Consider a system with $N$ sites per unit cell. Without loss of generality, we may assume that $N$ is even. We label the amplitude in the $m^{\text{th}}$ site as $a_{m}$. In general, due to $PT$ symmetry, the Bloch eigenvector will satisfy
@@
$\left|K\right\rangle=\frac{\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}}{\sqrt{1+r_{2}^{2}(K)+\dots}}\begin{pmatrix}\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu\phi(K)/2}\\ r_{2}(K)\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu\phi_{2}(K)}\\ \vdots\\ r_{2}(K)\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi_{2}(K)}\\ \mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}\end{pmatrix}.$ (31)
Note that the choice of reciprocal gauge $\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\phi(K)/2}$ is physical: It is required for the Zak phase to have an unambiguous interpretation of mean displacement. To prove this, we first fix the first site to have position $x_{1}=0$, so that the lattice position of the $m^{\text{th}}$ site is given by $x_{m}=m-1$. As a result, the Bloch eigenvector is related to the discrete fourier transform (DFT) by
@@
$\langle m=1|K\rangle=\langle m=1|K,\text{DFT}\rangle=a_{1}=\text{const.}$. (34)
We may fix the global phase to 00, so that the Bloch eigenvector is given by (31). We thus see that allowing the phase of the first entry in $\left|K\right\rangle$ to vary is equivalent to choosing a different, translated coordinate system for each value of the system quasi momentum $K$.
@@
$=\left\langle K,\text{Sym}\right|\left[-\mathrm{i}\mkern 1.0mu\partial_{K}+\frac{\phi^{\prime}(K)}{2}\right]\left|K,\text{Sym}\right\rangle.$ (35)
While computing the inner product $\left\langle K,\text{Sym}\right|\partial_{K}\left|K,\text{Sym}\right\rangle$, the $m^{\text{th}}$ and $(N-m)^{\text{th}}$ terms always cancel, so we are left with
@@
The generalisation to continuous systems is straightforward by taking $N\to\infty$ with a symmetric finite difference scheme.
Recently, it has been brought to our attention that [40] has previously obtained a general expression for the Zak phase in terms of the eigenvalue of the inversion operator acting on $\left|K=0\right\rangle$ and $\left|K=\pi\right\rangle$, which can be used to derive Eq. (12) from the main text more straightforwardly.
Appendix B Four Branches of Solutions for $T$
In this appendix section, we clarify the four branches of solutions mentioned in Sec. III.2. From parity symmetry, we have the constraint given in (10). Substituting this back into $T$ and taking the trace gives us
$\mathrm{Tr}(T)=\sec(k)\left(a\cos(2k)+\frac{1}{a}\right)$ (36)
Solving for the condition $|\mathrm{Tr}(T)|=2$ then, we obtain four different solutions
$a=\frac{1}{\pm\cos(k)\pm\sin(k)}\quad\text{or}\quad\frac{1}{\pm\cos(k)\mp\sin(k)}.$ (37)
Since $a>0$ we are required to choose only the positive solutions in Eq. (37) for each $k$. Substituting (37) back into $T$ gives us the four branches mentioned in the main text. Note that the conditions $\det(T)=1$ and $|\operatorname{Tr}(T)|=|2\cos K|=2$ fix all eigenvalues to have the same sign, and be ±$\pm 1$. This way, $K$ is quantized to be $K\in\pi\mathbb{Z}$. In the case $k\not\in\pi\mathbb{Z}$ these solutions will not be diagonalisable and the Bloch state can be recovered from the unique $T$-eigenvector. The results are summarised in Table 1.
Table 1: Table summarising the four solutions of $T\in\mathrm{SL}(2;\mathbb{R})$ satisfying both (10) and $|\mathrm{Tr}(T)|=2$. The value of $K$ is obtained as the argument of the eigenvalue of $T$. The value of $\varphi$ is the phase difference between the two entries of the unique eigenvector of $T$.
$T$ $K$ $\varphi$
$\begin{pmatrix}-\frac{\cos(k)}{\sin(k)+\cos(k)}&\frac{\sin(k)}{\sin(k)+\cos(k)}\\ -\frac{\sin(k)}{\sin(k)+\cos(k)}&-\frac{2\sin(k)+\cos(k)}{\sin(k)+\cos(k)}\end{pmatrix}$ $\pi$ $\pi$
$\left(\begin{array}[]{cc}\frac{\cos(k)}{\cos(k)-\sin(k)}&-\frac{\sin(k)}{\cos(k)-\sin(k)}\\ \frac{\sin(k)}{\cos(k)-\sin(k)}&\frac{\cos(k)-2\sin(k)}{\cos(k)-\sin(k)}\end{array}\right)$ 0 0
$\left(\begin{array}[]{cc}\frac{\cos(k)}{\sin(k)-\cos(k)}&\frac{\sin(k)}{\cos(k)-\sin(k)}\\ \frac{\sin(k)}{\sin(k)-\cos(k)}&-\frac{\cos(k)-2\sin(k)}{\cos(k)-\sin(k)}\end{array}\right)$ $\pi$ 0
$\left(\begin{array}[]{cc}\frac{\cos(k)}{\sin(k)+\cos(k)}&-\frac{\sin(k)}{\sin(k)+\cos(k)}\\ \frac{\sin(k)}{\sin(k)+\cos(k)}&\frac{2\sin(k)+\cos(k)}{\sin(k)+\cos(k)}\end{array}\right)$ 0 $\pi$
Appendix C Coupled Mode Approximation to the Dimerised Photonic Crystal